# TerraFusion: Joint Generation of Terrain Geometry and Texture Using Latent Diffusion Models

### v1

@@ 1 Introduction
Compared to baselines, our approach produces better qualitative and quantitative results and more accurately captures the correlation between heightmaps and textures.
@@ 4.1 Dataset
We used NASADEM elevation data [^1] https://lpdaac.usgs.gov/products/nasadem_hgtv001/ (30 m spatial resolution) for heightmaps, and atmospherically corrected Sentinel-2 imagery [^2] https://sentiwiki.copernicus.eu/web/s2-processing#S2Processing-CopernicusSentinel-2Collection-1AvailabilityStatus (Level-2A product, 10 m spatial resolution) for terrain textures. We focused on natural terrain between 60°S and 60°N, extracting $\tcdegree\times 1\tcdegree$ 1 geographic regions from 276 land areas (see Appendix A for the details). Each region was resampled to a spatial resolution of 25 m, producing images of approximately $4096\times 4096$ pixels. These were divided into non-overlapping $256\times 256$ patches, excluding data with missing texture details due to clouds and shadows. To match the input size required by the diffusion model, each patch was then upsampled to $512\times 512$. The elevation values in the heightmaps ranged from 0 to 8000 m. However, as discussed in Section 4.2, VAEs struggle to capture this wide distribution accurately. To mitigate this, we excluded minor samples with elevations above 2000 m. This preprocessing yielded a total of 4,119 paired samples of heightmaps and texture images.
@@ 4.2 VAE Training
It is a standard practice to normalize the inputs and outputs of a VAE to the range $[-1,1]$ for compatibility with activation functions. Accordingly, we normalize heightmaps in this study using the following equation:
@@ Appendix A Detail of Terrain Dataset
For texture acquisition, we targeted land areas between 60°S and 60°N observable by Sentinel-2. We randomly sampled 500 locations, each covering a $\tcdegree\times 1\tcdegree$ 1 region in latitude and longitude. To focus on natural terrain, we excluded regions with an average elevation below 100 meters, which are likely to include urban or artificial structures. We further filtered locations using the Global Human Modification dataset [11], which quantifies human impact on the environment (spatial resolution: 1000 m; value range: [0, 1], with higher values indicating greater modification). Regions with a human modification index of 0.3 or higher were excluded. To mitigate climate-related sampling bias, we applied constraints based on the Köppen climate classification. We excluded arid zones and focused on four climate categories: tropical, temperate, subarctic, and polar. Sampling was limited to a maximum of 125 locations per category to maintain climate balance. We did not enforce uniform sampling (e.g., 100 per category), as some climate zones contained fewer than 100 suitable regions. After sampling, we applied additional filtering using the ESA WorldCover dataset [22] (spatial resolution: 10 m). We removed images including areas labeled as cropland, built-up regions, or water bodies, as were regions obscured by clouds or cloud shadows. As a result, we obtained usable data from 276 locations.

### v2

@@ 1 Introduction
Compared to baselines, our approach produces better qualitative and quantitative results and more accurately captures the correlation between heightmaps and textures. Our code and pretrained models are available at https://github.com/millennium-nova/terra-fusion.git.
@@ 4.1 Dataset
We used NASADEM elevation data [^1] https://lpdaac.usgs.gov/products/nasadem_hgtv001/ (30 m spatial resolution) for heightmaps, and atmospherically corrected Sentinel-2 imagery [^2] https://sentiwiki.copernicus.eu/web/s2-processing#S2Processing-CopernicusSentinel-2Collection-1AvailabilityStatus (Level-2A product, 10 m spatial resolution) for terrain textures. We focused on natural terrain between 60°S and 60°N, extracting $1\tcdegree\times 1\tcdegree$ geographic regions from 276 land areas (see Appendix A for the details). Each region was resampled to a spatial resolution of 25 m, producing images of approximately $4096\times 4096$ pixels. These were divided into non-overlapping $256\times 256$ patches, excluding data with missing texture details due to clouds and shadows. To match the input size required by the diffusion model, each patch was then upsampled to $512\times 512$. The elevation values in the heightmaps ranged from 0 to 8000 m. However, as discussed in Section 4.2, VAEs struggle to capture this wide distribution accurately. To mitigate this, we excluded minor samples with elevations above 2000 m. This preprocessing yielded a total of 4,119 paired samples of heightmaps and texture images.
@@ 4.2 VAE Training
It is a standard practice to normalize the inputs and outputs of a VAE to the range [1,1][-1,1] for compatibility with activation functions. Accordingly, we normalize heightmaps in this study using the following equation:
@@ 5.1 Quantitative Evaluation
Table 3: Statistics of correlation coefficients ($\uparrow$) between textures and heightmaps generated by the two-stage baselines and our method. IQR denotes the interquartile range. Values in parentheses indicate absolute differences from the ground-truth (training data). The best scores are shown in bold, and the second-best are underlined.
Method mean std 25% 50% 75% IQR
Baseline ($\mathbf{h}\rightarrow\mathbf{x}$) 0.13 (0.06) 0.11 (0.03) 0.05 (0.02) 0.10 (0.06) 0.18 (0.10) 0.13 (0.07)
Baseline ($\mathbf{x}\rightarrow\mathbf{h}$) 0.12 (0.07) 0.10 (0.04) 0.04 (0.03) 0.10 (0.06) 0.18 (0.10) 0.13 (0.07)
Ours (from scratch) 0.17 (0.02) 0.12 (0.02) 0.07 (0.00) 0.15 (0.01) 0.25 (0.03) 0.18 (0.02)
Ours (fine-tuning) 0.16 (0.03) 0.12 (0.02) 0.06 (0.01) 0.13 (0.03) 0.23 (0.05) 0.17 (0.03)
GT (train data) 0.19 0.14 0.07 0.16 0.28 0.20
Table 4: Statistics of correlation coefficients ($\uparrow$) between textures and heightmaps generated by PSGAN [8] and our method. IQR denotes the interquartile range. Values in parentheses indicate the absolute differences from the GT (training data)
Method mean std 25% 50% 75% IQR
PSGAN 0.15 (0.04) 0.11 (0.03) 0.06 (0.01) 0.13 (0.03) 0.22 (0.06) 0.16 (0.04)
Ours (from scratch) 0.17 (0.02) 0.12 (0.02) 0.07 (0.00) 0.15 (0.01) 0.25 (0.03) 0.18 (0.02)
Ours (fine-tuning) 0.16 (0.03) 0.12 (0.02) 0.06 (0.01) 0.13 (0.03) 0.23 (0.05) 0.17 (0.03)
GT (train data) 0.19 0.14 0.07 0.16 0.28 0.20
@@
We also evaluate how well each model captures the correlation between heightmaps and textures. For each sample, we compute the Pearson correlation coefficient between the heightmap and each texture channel, and then average across channels. The distributions of these correlation coefficients are visualized using violin plots. A distribution closer to that of the training data indicates better modeling of real-world terrain relationships. Figure 8 presents the results. Overall, the two-stage models produce lower correlation values compared to the training data, while our models yield distributions that more closely match the training distribution. Table 3 summarizes the differences in correlation values between each method and the ground-truth data, further demonstrating the effectiveness of the joint generative models. In addition, Figure 9 and Table 4 show comparisons between our method and PSGAN [2]. Although the correlation values of the two methods are comparable, PSGAN often produces artifacts in heightmaps, as discussed in Section 5.2.
@@ 5.2 Qualitative Evaluation
Texture Conditioning.
To enable more flexible terrain generation, we explored texture control. As shown on the left side of Figure 16, we first created a two-color-quantized texture dataset by approximating each texture image in the satellite image dataset with two representative colors. ControlNet was then trained using this dataset, following the same procedure described in Figure 7. Given a user-drawn two-color image, the model synthesizes a texture conditioned on that image. Representative results are shown on the right side of Figure 16. These textures reflect the user-drawn inputs; however, the current model struggles to capture global terrain features, indicating room for improvement. Producing terrain-like outputs also requires complex input cues, and enabling simpler inputs remains an open challenge.
Figure 16: Details on our experiment on texture conditioning using ControlNet.
6 Discussion
We demonstrate that our method outperforms GAN-based approaches [2, 22], which represent the state of the art in unconditional terrain generation. A recent concurrent study, MESA [3], employs a diffusion-based model for terrain synthesis, as discussed in Section 2. However, MESA focuses on text-driven terrain generation, which differs from the unconditional and sketch-based tasks addressed in our work, making direct comparison beyond the scope of this study. We regard our method as complementary, and extending it to support textual input remains an interesting direction for future research.
Evaluation by professional artists.
We asked professional artists at the game studio Polyphony Digital Inc. to review our generated terrains and provide feedback. They noted that the textures vary plausibly with local slope and elevation, suggesting that our framework successfully captures essential geometry–appearance correlations. However, they also identified several limitations. Some textures contain baked-in shading, indicating the need for shading-free (albedo) textures via de-lighting. In addition, low-frequency noise sometimes causes blurring in the textures. In terms of geometry, the relatively low resolution of both heightmaps and textures limited the evaluation of fine-scale realism relative to real-world terrains. The generated results also lacked key geomorphological features such as rivers, lakes, tree lines, and snow cover, which are important for producing convincing landscapes. The artists further emphasized the importance of explicitly defining spatial scale (e.g., $500~m\times 500~m$ versus $10~km\times 10~km$), because recognizable landforms depend on scale. For more in-depth evaluation, they suggested focusing on specific geographic regions to simplify the problem, improve output quality, and assess results in the context of region-specific landforms.
Resolution
Resolution. Although our current experiments are conducted at a resolution of 512×512, scaling diffusion-based terrain generation to higher resolutions remains a significant challenge. While GANs are often considered efficient for high-resolution synthesis, recent advances in latent diffusion models (LDMs), such as SDXL [19], demonstrate comparable capabilities. We believe our framework can be extended to higher resolutions by leveraging such improved LDMs, as well as approaches like latent-space super-resolution [10], which support coherent upsampling in the latent domain and help preserve structural fidelity and fine detail. Furthermore, methods such as consistency models [21], which reduce the number of denoising steps, may lower computational costs during high-resolution inference. Exploring these approaches is an important direction for future work.
@@ Appendix A Detail of Terrain Dataset
For texture acquisition, we targeted land areas between 60°S and 60°N observable by Sentinel-2. We randomly sampled 500 locations, each covering a $1\tcdegree\times 1\tcdegree$ region in latitude and longitude. To focus on natural terrain, we excluded regions with an average elevation below 100 meters, which are likely to include urban or artificial structures. We further filtered locations using the Global Human Modification dataset [12], which quantifies human impact on the environment (spatial resolution: 1000 m; value range: [0, 1], with higher values indicating greater modification). Regions with a human modification index of 0.3 or higher were excluded. To mitigate climate-related sampling bias, we applied constraints based on the Köppen climate classification. We excluded arid zones and focused on four climate categories: tropical, temperate, subarctic, and polar. Sampling was limited to a maximum of 125 locations per category to maintain climate balance. We did not enforce uniform sampling (e.g., 100 per category), as some climate zones contained fewer than 100 suitable regions. After sampling, we applied additional filtering using the ESA WorldCover dataset [25] (spatial resolution: 10 m). We removed images including areas labeled as cropland, built-up regions, or water bodies, as were regions obscured by clouds or cloud shadows. As a result, we obtained usable data from 276 locations.